OpenStudy (anonymous):

2/3 of 25% of 1/2 of 2000

OpenStudy (anonymous):

Anyone

OpenStudy (math&ing001):

\[\frac{ 2 }{ 3 }*\frac{ 25 }{ 100 }*\frac{ 1 }{ 2 }*2000\]